OpenStudy (anonymous):

\[\huge 2^{2x+2}-a*2^{x+2}+5-4a \ge 0 \forall x \epsilon R \] is true. then find "a".

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

Let: \[\large 2^{x+2} = y\] \[y^2 - ay + (5-4a) \ge 5\]

OpenStudy (anonymous):

No, the question is wrong..

OpenStudy (anonymous):

\[y^2 - ay + (5-4a) \ge 0\]

OpenStudy (anonymous):

oh sorry question is right and @waterineyes y^2/4 -ay+(5-4a) ..... ------

OpenStudy (anonymous):

If the question is right then you are also right.. @mukushla benim yoldashim..

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

Then we will get: \[y^2 - 4ay + (20 - 16a) \ge 0\]

OpenStudy (anonymous):

Find Discriminant from here..

OpenStudy (anonymous):

Or I am wrong???

OpenStudy (anonymous):

quite right

OpenStudy (anonymous):

\(D = b^2 - 4ac\) \(D = 16a^2 - 80 + 64a\) \(D = 80a - 80\)

OpenStudy (anonymous):

D=16(a^2+4a-5) be carefull

OpenStudy (anonymous):

Oh sorry sorry..

OpenStudy (anonymous):

\[D = 16a^2 - 80 + 64a\] Take 16 common: \[D = 16(a^2 +4a -5)\]

OpenStudy (anonymous):

We have to solve this equation or not??

OpenStudy (anonymous):

\[D = 16(a+5)(a-1)\]

OpenStudy (anonymous):

we have to solve \( D \le 0 \)

OpenStudy (anonymous):

Yes.. \[(a+5)(a-1) \le 0\]

OpenStudy (anonymous):

a = 1 and a = -5 is it??

OpenStudy (anonymous):

Or, a < 1 and a > -5??

OpenStudy (anonymous):

\[-5 \le a \le 1\]